an electrol cell is composed of Cu and Zn.A current of 9.65A is drawn from a cell for 1 HR.THEN THE LOSS IN mass a anode and gain in cathode ,respevct could be.?

Grade:12

1 Answers

Naveen Kumar
9 years ago
The cell reaction can be written as:
Cu2+.....+.......Zn.................................>Cu............+........Zn2+
so for this process if 1 mole of Cu is deposited, moles of Zn reacted=1mole and the moles of electron involved=2mole=2F charge
Now calculkate the charge supplied in 1Hr,
Q=i*t=9.65*60*60 C
To calculate the charge on 1 mole of electron, we should multiply 1.6*10^(-19)C * 6.023*10^23=1F
so convert Q from C to F
now calculate moles of electron in this Q charge and then for each one mole electron ½ mole of Cu would be deposited and hence weight increase on cathode=moles of Cu*(mol. weight of Cu)
Similarly for the one mole electron thethe weight increase =1/2*(moles of Zn)